# TSkips: Efficiency Through Explicit Temporal Delay Connections in Spiking Neural Networks

### v1

@@ 8.2 Convergence Analysis
8.3 TSkips in ANNs and Hybrid Models
@@
Table 7: Classification accuracy of ANN and Hybrid ANN-SNN models on the SHD and SSC datasets. We compare the 4-layer MLP baseline (Baseline-1) with and without Forward (F) and Backward (B) TSkips.
@@
F TSkip 0.24 69.08 84.54
B TSkip 0.19 73.45 86.02
F + B TSkip 0.20 73.25 85.74
@@
F TSkip 0.42 71.19 73.48
B TSkip 0.24 70.23 72.87
F + B TSkip 0.54 71.09 74.96
@@
8.4 Comparison between TSkip Operators
To further analyze the impact of different TSkip connection types, we compared the performance of concatenation-based and addition-based TSkips on the 4-layer and 8-layer baselines (Table 3). Table 8 demonstrates the superior performance of concatenation-based TSkips over addition-based on the baseline models. Concatenation-based TSkips consistently outperformed their addition-based counterparts. This can be attributed to the increased representational capacity of concatenation, which expands the feature space by combining information from the skip connection and the destination layer. This richer representation allows the network to learn more complex temporal patterns, particularly crucial in deep SNNs where information is encoded in sparse spike trains and membrane potentials. Furthermore, concatenation preserves information from both sources, preventing potential information loss that can occur with addition.
Table 8: Comparison of classification accuracy on SHD and SSC datasets with concatenation-based (C) and addition-based (A) Forward (F) and Backward (B) TSkips on a 4-layer (Baseline -1) and 8-layer (Baseline - 2) MLP network.
@@ 8.4 Comparison between TSkip Operators
F TSkip - C 0.24 0.42 92.32 76.5
B TSkip - C 0.19 0.24 93.64 79.87
F + B TSkip - C 0.20 0.54 93.01 78.64
F TSkip - A 0.19 0.37 87.65 72.13
B TSkip - A 0.11 0.21 87.54 73.15
F + B TSkip - A 0.11 0.40 88.13 75.12
@@
F TSkip - C 1.12 1.08 94.15 78.98
B TSkip - C 1.16 1.14 93.86 79.65
F + B TSkip - C 1.28 1.42 94.73 80.23
F TSkip - A 0.99 1.05 86.39 72.15
B TSkip - A 0.97 1.01 87.32 75.53
F + B TSkip - A 0.98 1.14 89.45 77.12
8.5 Hyper parameters for DVS128 Gesture, SHD and SSC
Table 9 presents the top-ranked network architectures and TSkip configurations identified by [27], with corresponding results shown in Tables  2 and 3.
@@ 8.5 Hyper parameters for DVS128 Gesture, SHD and SSC
      SHD and SSC: Specifies the number of input channels for each MLP layer, including the final fully connected layer. Channels represented as $n*2$ indicate that concatenation-based TSkips were used, doubling the number of channels at that layer. This representation is used to highlight that the results shown in Table 8, which uses addition-based TSkips, are based on the same underlying network architectures.
@@
    TSkip parameters: The temporal delay ($\Delta t$) and the origin and destination layers $(from\rightarrow to)$ for Forward (F), Backward (B), and combined (F+B) TSkips.
@@
      Leak and threshold initialization values for the adaptive LIF neurons: 0.6 and 15, respectively.
@@
Variant Network Delay ($\Delta t)$ TSkip Position (from $\rightarrow$ to)
DVS128 Gesture
Baseline 2$\times$64$\times$64-3c80s1-3c80s1- 5c86s1-5c64s1-1c64s1-1c32s11 - -
F TSkip 2$\times$64$\times$64-3c83s1-3c83s1-5c110s1- 3c94s1-1c94s1-1c32s11 5 1 $\rightarrow$ 2
B TSkip 2$\times$64$\times$64-3c64s1-5c64s1- 5c64s1-5c76s1-1c76s1-1c32s11 5 4 $\rightarrow$ 3
F + B TSkip 2$\times$64$\times$64-3c122s1-5c122s1- 3c122s1-5c64s1-1c64s1+1c32s11 F: 6 B: 8 F: 1 $\rightarrow$ 3 B: 5 $\rightarrow$ 4
SHD
Baseline - 1 700-124-288-144-20 - -
F TSkip - 1 700-124*2-432-115-20 24 1 $\rightarrow$ 2
B TSkip - 1 700*2-124-115-96-20 14 4 $\rightarrow$ 1
F + B TSkip - 1 700*2-124-144-72*2-20 F: 18 B: 14 F: 2 $\rightarrow$ 4 B: 2 $\rightarrow$ 1
Baseline - 2 700-512-288-224- 192-896-288-20 - -
F TSkip - 2 700-532-288-524- 192-448*2-288-20 17 2 $\rightarrow$ 6
B TSkip - 2 700-321-558*2-334 292-452-328-20) 12 3 $\rightarrow$ 5
F + B TSkip - 2 700-321-228*2-334- 492*2-462-448-20 F: 10 B: 13 F: 2 $\rightarrow$ 5 B: 7 $\rightarrow$ 3
SSC
Baseline - 1 700-124-148-130-35 - -
F TSkip - 1 700-324-298*2-145-35 15 1 $\rightarrow$ 3
B TSkip - 1 700-184*2-179-230-35 12 3 $\rightarrow$ 2
F + B TSkip - 1 700-364*2-257*2-195-35 F: 14 B: 8 F: 1 $\rightarrow$ 3 B: 4 $\rightarrow$ 2
Baseline - 2 700-731-410-340- 150-184-58-35 - -
F TSkip - 2 700-646-410-400- 250-184*2-135-35 17 2 $\rightarrow$ 6
B TSkip - 2 700-663-373*2-337- 292-182-135-35 15 7 $\rightarrow$ 3
F + B TSkip - 2 700-674-493*2-347*2- 278-192-89-35 F: 9 B: 11 F: 1 $\rightarrow$ 2 B: 7 $\rightarrow$ 3
9 NASWOT-SAHD Correlation on TSkips
To validate the effectiveness of NASWOT-SAHD [27] in identifying optimal network configurations, we analyzed the correlation between the accuracy and score of the identified networks, as shown in Fig. 11. The scores are based on the Sparsity Aware Hamming Distance (SAHD) [27], and the correlation is measured using Kendall’s $\tau$. Our baseline models achieved a $\tau$ correlation of 0.63, consistent with the results reported in [27]. However, the forward and backward TSkip network searches resulted in slightly lower correlations of 0.58 and 0.55, respectively. All results reported are for the top-ranked networks identified by  [27].
@@ 9 NASWOT-SAHD Correlation on TSkips
    TSkip Position: We constrained connections to not be within the same layer, ensuring that all connections originated from a different layer than the destination.
@@
      DSEC-Flow: No constraint was applied as we directly replaced existing skip connections in the fully spiking[29] and hybrid [36] EV-FlowNet architecture with TSkips, maintaining the original model size.
@@
Figure 11: Correlation between accuracy and SAHD score for different network configurations (baseline, forward TSkip, and backward TSkip).

### v2

@@ C.3 Convergence Analysis
C.4 TSkips in ANNs and Hybrid Models
@@
Table 10: Classification accuracy of ANN and Hybrid ANN-SNN models on the SHD and SSC datasets. We compare the 4-layer MLP baseline (Baseline-1) with forward/backward TSkips.
@@
Forward TSkips 0.24 69.08 84.54
Backward TSkips 0.19 73.45 86.02
Forward + Backward TSkips 0.20 73.25 85.74
@@
Forward TSkips 0.42 71.19 73.48
Backward TSkips 0.24 70.23 72.87
Forward + Backward TSkips 0.54 71.09 74.96
@@
C.5 Comparison between TSkip Operators
To further analyze the impact of different TSkips connection types, we compared the performance of concatenation-based and addition-based TSkips on the 4-layer and 8-layer baselines (Table 3).
Table 11: Comparison of classification accuracy on SHD and SSC datasets with concatenation-based (C) and addition-based (A) forward/backward TSkips on a 4-layer (Baseline -1) and 8-layer (Baseline - 2) MLP network.
@@ C.5 Comparison between TSkip Operators
Forward TSkips - C 0.24 0.42 92.32 76.5
Backward TSkips - C 0.19 0.24 93.64 79.87
Forward + Backward TSkips - C 0.20 0.54 93.01 78.64
Forward TSkips - A 0.19 0.37 87.65 72.13
Backward TSkips - A 0.11 0.21 87.54 73.15
Forward + Backward TSkips - A 0.11 0.40 88.13 75.12
@@
Forward TSkips - C 1.12 1.08 94.15 78.98
Backward TSkips - C 1.16 1.14 93.86 79.65
Forward + Backward TSkips - C 1.28 1.42 94.73 80.23
Forward TSkips - A 0.99 1.05 86.39 72.15
Backward TSkips - A 0.97 1.01 87.32 75.53
Forward + Backward TSkips - A 0.98 1.14 89.45 77.12
Table 11 demonstrates the superior performance of concatenation-based TSkips over addition-based on the baseline models. Concatenation-based TSkips consistently outperformed their addition-based counterparts. This can be attributed to the increased representational capacity of concatenation, which expands the feature space by combining information from the skip connection and the destination layer. This richer representation allows the network to learn more complex temporal patterns, particularly crucial in deep SNNs where information is encoded in sparse spike trains and membrane potentials. Furthermore, concatenation preserves information from both sources, preventing potential information loss that can occur with addition.
C.6 Hyper parameters for DVS128 Gesture, SHD and SSC
Table 13 presents the top-ranked network architectures and TSkips configurations identified by NASWOT-SAHD (Kim et al., 2022), with corresponding results shown in Tables  2 and 3.
@@ C.6 Hyper parameters for DVS128 Gesture, SHD and SSC
      SHD and SSC: Specifies the number of input channels for each MLP layer, including the final fully connected layer. The channels represented as $n*2$ indicate that concatenation-based TSkips were used, doubling the number of channels $(n)$ in that layer. This representation is used to highlight that the results shown in Table 11, which uses addition-based TSkips, are based on the same underlying network architectures.
@@
    TSkips parameters: The temporal delay ($\Delta t$) and the origin and destination layers $(from\rightarrow to)$ for forward (F), backward (B) and combined (F + B) TSkips.
@@
      Leak and threshold initialization values for adaptive LIF neurons: 0.6 and 15, respectively.
@@
Appendix D NASWOT-SAHD Validation
D.1 Correlation on TSkips
Figure 12: Correlation between accuracy and SAHD score for different network configurations (baseline, forward TSkips, and backward TSkips).
To validate the effectiveness of NASWOT-SAHD (Kim et al., 2022) in identifying optimal network configurations, we analyzed the correlation between the accuracy and score of the identified networks, as shown in Fig. 12. The scores are based on the Sparsity Aware Hamming Distance (SAHD) (Kim et al., 2022), and the correlation is measured using Kendall’s $\tau$. Our baseline models achieved a $\tau$ correlation of 0.63, consistent with the results reported in Kim et al. (2022). However, the forward and backward TSkips network searches resulted in slightly lower correlations of 0.58 and 0.55, respectively. All results reported are for the top-ranked networks identified by NASWOT-SAHD  (Kim et al., 2022).
@@ D.1 Correlation on TSkips
    TSkips Position: We constrained connections to not be within the same layer, ensuring that all connections originated from a different layer than the destination.
@@
      DSEC-Flow: No constraint was applied as we directly replaced existing skip connections in the fully spiking(Kosta & Roy, 2023) and hybrid (Negi et al., 2024) EV-FlowNet architecture with TSkips, maintaining the original model size.
@@
D.2 Robustness across Random Seeds
To validate the robustness of our NAS-identified configurations and the consistent performance enhancement provided by TSkips, we conducted additional experiments on the DSEC-flow (Gehrig et al., 2021a) dataset. Table 12 presents the Average Endpoint Error (AEE) for both baseline models and their TSkips-augmented variants, averaged across three different random seeds. These results, reported as mean $(\mu)\pm$ standard deviation $(\sigma)$, demonstrate the stability and reliable improvement offered by our approach.
Table 12: DSEC flow results on performing NAS search for TSkips with multiple random seeds. Presented below is the AEE (mean$(\mu)$$\pm$ standard deviation$(\sigma)$) (lower is better) of the models found across 3 random seeds.
Model $\Delta t$ Position AEE ($\mu\pm\sigma$)
Base - Baseline - - 1.38 $\pm$ 0.09
Base - FTSkip 3,2,3 1,1,2 1.16 $\pm$ 0.10
Base - BTSkip 4,3,4 1,2,1 1.14 $\pm$ 0.13
Mini - Baseline - - 1.72 $\pm$ 0.07
Mini - FTSkip 3,2,2 1,1,2 1.49 $\pm$ 0.06
Mini - BTSkip 3,2,3 2,1,2 1.47 $\pm$ 0.07
Micro - Baseline - - 1.97 $\pm$ 0.11
Micro - FTSkip 4,3,3 2,1,2 1.66 $\pm$ 0.10
Micro - BTSkip 3,2,2 3,2,3 1.69 $\pm$ 0.15
Nano - Baseline - - 2.23 $\pm$ 0.17
Nano - FTSkip 3,3,2 1,1,2 1.85 $\pm$ 0.07
Nano - BTSkip 4,3,3 2,1,2 1.78 $\pm$ 0.13
Pico - Baseline - - 2.64 $\pm$ 0.08
Pico - FTSkip 4,2,3 1,2,1 2.32 $\pm$ 0.13
Pico - BTSkip 3,2,3 2,1,1 2.28 $\pm$ 0.17
Variant Network
Delay
($\Delta t$)
TSkip
Position
(from $\rightarrow$ to)
DVS128 Gesture
Baseline
2$\times$64$\times$64-3c80s1-3c80s1-
5c86s1-5c64s1-1c64s1-1c32s11
F TSkips
2$\times$64$\times$64-3c83s1-3c83s1-5c110s1-
3c94s1-1c94s1-1c32s11
5 1 $\rightarrow$ 2
B TSkips
2$\times$64$\times$64-3c64s1-5c64s1-
5c64s1-5c76s1-1c76s1-1c32s11
5 4 $\rightarrow$ 3
F + B
TSkips
2$\times$64$\times$64-3c122s1-5c122s1-
3c122s1-5c64s1-1c64s1+1c32s11
F: 6
B: 8
F: 1 $\rightarrow$ 3
B: 5 $\rightarrow$ 4
SHD
Baseline - 1 700-124-288-144-20 - -
F TSkips - 1 700-124*2-432-115-20 24 1 $\rightarrow$ 2
B TSkips - 1 700*2-124-115-96-20 14 4 $\rightarrow$ 1
F + B
TSkips
- 1
700*2-124-144-72*2-20
F: 18
B: 14
F: 2 $\rightarrow$ 4
B: 2 $\rightarrow$ 1
Baseline - 2
700-512-288-224-
192-896-288-20
F TSkips - 2
700-532-288-524-
192-448*2-288-20
17 2 $\rightarrow$ 6
B TSkips - 2
700-321-558*2-334
292-452-328-20)
12 3 $\rightarrow$ 5
F + B
TSkips
- 2
700-321-228*2-334-
492*2-462-448-20
F: 10
B: 13
F: 2 $\rightarrow$ 5
B: 7 $\rightarrow$ 3
SSC
Baseline - 1 700-124-148-130-35 - -
F TSkips - 1 700-324-298*2-145-35 15 1 $\rightarrow$ 3
B TSkips - 1 700-184*2-179-230-35 12 3 $\rightarrow$ 2
F + B
TSkips
- 1
700-364*2-257*2-195-35
F: 14
B: 8
F: 1 $\rightarrow$ 3
B: 4 $\rightarrow$ 2
Baseline - 2
700-731-410-340-
150-184-58-35
F TSkips - 2
700-646-410-400-
250-184*2-135-35
17 2 $\rightarrow$ 6
B TSkips - 2
700-663-373*2-337-
292-182-135-35
15 7 $\rightarrow$ 3
F + B
TSkips
- 2
700-674-493*2-347*2-
278-192-89-35
F: 9
B: 11
F: 1 $\rightarrow$ 2
B: 7 $\rightarrow$ 3